### v3

@@ 1 Introduction
Landslides are severe geological hazards that often results in traffic disruption, village burial, river blockage and other catastrophic incidents, posing significant threats to human life and property. For instance, Lanzhou City, China, has experienced 24 large-scale landslides since 1949, resulting in 670 fatalities and direct economic losses of 776 million RMB [34]. Therefore, accurate and efficient landslide prediction is crucial for preventing and reducing such disastrous losses. Besides, studying existing landslides could provide valuable insights for predicting potential landslides that lack fully-formed features and present the greatest risks, so this paper focuses on improving the detection of existing landslides.
@@
Traditional methods for landslide detection rely on field investigations and remote sensing data, which, though reliable, are time-consuming, inefficient, and highly dependent on experts experience, therefore hard to be widely applied. With the development of remote sensing technologies, abundant ground-observable data has been introduced for automatic landslide detection. For example, Interferometric Synthetic Aperture Radar (InSAR) data can provide deformation characteristics [26, 46], High-Resolution Satellite Image (HRSI) data can offer optical features [28, 27], and Digital Elevation Model (DEM) and Digital Surface Model (DSM) data can present topographic information [20, 3]. There are two classic methods using HRSI: pixel-based classification and object-oriented analysis (OBA). Pixel-based methods classify pixels based on spectral features [25, 45, 22] but suffer from salt-and-pepper noise and lack the ability to identify spatially continuous regions [38, 35], whereas OBA methods improve the spatial continuity by grouping adjacent pixels into segments or objects with a homogeneity factor, leveraging spatial, texture, context, geometry, and spectral features [1, 29, 24, 16, 32]. However, OBA methods require manual adjustments for thresholds such as relief and slopel, limiting their scalability.
@@
Recent machine learning methods including support vector machine, random forest, logistic regression [4, 10, 31], and convolutional neural networks (CNNs) have demonstrated improved accuracy and efficiency in landslide detection tasks. CNNs in particular are widely applied due to their excellent end-to-end self-learning ability and abstract representation capability. In [12] a Siamese CNN based on dual-temporal landslide HRSI data is developed for the pixel-wise change detection of landslides and a generative adversarial network is introduced to suppress differences between two images in the time domain. In [19], a 3D spatial and channel attention module is designed to enhance the extraction of landslide features. In [41], a dual path attention network is proposed to leverage both texture and spatial information in remote sensing images.
@@
Figure 1: An example to show the visual differences between an relic landslide and a new landslide: a relic landslide, b new landslide
@@
Despite significant progress, most research focuses on detecting new landslides [12, 14, 19, 21, 2, 41] rather than relic landslides [44, 11, 27, 28]. Since landslide morphology gradually changes over time, there is a significant difference in morphology between new and relic landslides. New landslides, in a state of repeated or just stopped activity, present visually obvious features, such as distinct colors and contours against background. Relic landslides, however, have been inactive for a long time and are severely affected by vegetation coverage, natural erosion, and human activities such as terracing, building houses, and constructing roads, exhibiting blurry visual features difficult to distinguish from background, as shown in Fig. 1. Therefore, relic landslides are much difficult to identify than new ones.
@@
To address this challenge, we propose a multi-modal data fusion method leveraging HRSI and DEM data, where DEM data can provide elevation features and gradient patterns that effectively complement HRSI’s optical information. The common data fusion strategies can be categorized into early fusion [14, 19] and late fusion [3, 33, 43, 39, 27]. The early fusion integrates multi-modal data at preprocessing stage such as combining several inputs from different modalities into one or mixing several datasets as a whole. For example, in [14], topographic images are regarded as additional channels of spectral images, and in [19], satellite optical images, shape files of landslides’ boundaries and DEM data are mixed into one integral dataset. The early fusion can provide richer information and is easy to implement, but it requires powerful models, large training dataset, and massive computational resources, making it unsuitable for scenarios involving small-sized datasets. Besides, due to the fact that different modalities of data often have different semantic and syntactic features (e.g., visual features such as morphology, texture, and color from HRSIs, and terrain features like height variation from DEM), the data space after fusion is always complex and irregular, thus making further feature extraction from it quite challenging. This calls for more powerful models which place higher demands on datasets. In contrast, the late fusion independently extract heterogeneous features and then fuse them in the semantic space, which effectively obtains the essential semantic information required for the task. For example, in [3], the semantic features from near infrared, red, green (IRRG) spectrum and Digital Surface Model (DSM) data are concatenated in the channel dimension; in [33], optical image features and DSM data features are contenated; in [43], two regional feature maps from optical image and Synthetic Aperture Radar data are contenated; in [27], the features from HRSI and DEM data are fused by adding pixels with the same positions in spatial dimension. Therefore, considering the heterogeneity between HRSIs and DEM data, we adopt the late fsuion. Instead of pixel-wise addition [27], we fuse by channel concat since it performs better when there are significant semantic differences between heterogeneous features, such as HRSI and DEM, whereas the pixel-wise addition increases less complexity but at the expense of the discriminative ability of the model for heterogeneous features.
@@
Figure 2: Illustration of projections from a 3D landslide (bottom) into DEM (left), Contour (middle) and HRSI (right). One of the salient features of a landslide is the steep altitude descent of back and side walls, which is located in light gray areas with high elevation in DEM. The contour map is obtained by applying the Gauss-Laplace edge operator to DEM, where the back wall is parallel to the contour lines, while the side wall is vertical to the contour lines highlighted in green. The yellow arrows indicate landslide aspect. The gradient variation features in DEM are projected as the texture features in HRSI. The bottom schematic diagram of landslide is taken from [13]
@@
In order to tackle the small-sized dataset problem, we employ contrastive learning to enhance feature extraction by distinguishing positive samples from negative ones in the feature space [42, 15, 8]. Based on expert knowledge that the boundary features of back and side walls reflect the most elevation change of landslides as illustrated in Fig. 2, we propose a supervised hyper-pixel-wise contrastive learning method forcing the model to learn these local salient features in the feature space with great efficiency and reliability. The hyper pixel, defined as an area of the original image corresponding to a pixel in the high-level feature map by position, contains much rich and abundant semantic information from a local region compared to the information captured by individual pixels in pixel-wise contrastive learning.
@@ 2.1 Study area
The study area locates in the Loess Plateau at the intersection of the western end of the Western Qinling Fold Belt and the southeastern end of the Qilian Fold Belt. Specifically, the relic landslides are distributed in six counties of Gansu Province of China, i.e., Kangle, Weiyuan, Lintan, Lintao, Guanghe, and Dangchang. Fig. 3 shows the location of the study area and the landslide distribution maps from Lintao and Weiyuan. And the distribution maps of the remain four counties are presented in Appendix (Fig. 14).
Figure 3: The geo-location of the study area and the distribution maps from Lintao and Weiyuan
The low mountain and hilly landforms in the area are generally developed, with vertical and horizontal gullies. The stratigraphic lithologies are dominated by loose loess on the surface and weak mudstone in the lower part, with residual and alluvial parent material of soil. Besides, most of the soil moisture conditions in the area are not suitable for forest development, and the natural vegetation is grassland and shrub. Coniferous forest and deciduous broad-leaved forest are distributed in southern part of the study area since climate there is relatively humid [47]. Therefore, the vegetation coverage and the viscous soil with weak water permeability, corrosion resistance and erosion resistance are attributed to the the occurrence of landslides in the past.
@@ 2.2 Loess Plateau relic landslide dataset
The dataset used in the paper were measured in 2018 by the Gaofen-1 satellite, and contains 168 relic landslide samples cropped to the size of $512\times 512$. The HRSIs are RGB images and the DEM data are grayscale images. The images in the dataset are divided into the training, validation, and test sets in a ratio of 6:2:2. The resolution of the HRSI is 2m/pixel, and the resolution of the DEM data is originally 30m/pixel and subsequently interpolated to match the resolution of the HRSI. Since there are only 266 landslide samples, 5-fold cross validation is employed to achieve more stable experimental results, which differs from the fixed dataset partition scheme in the reference baseline model [27]. Both the dataset partition scheme and the extended dataset cause the greater difference in numerical results of the baseline model.
@@ 3 Methods
Figure 4: Architecture of HPCL-Net
@@ 3.1 Encoder
As shown in Fig. 4, the encoder consists of HFE and MAFE. Improved from our previous work [27], the newly designed encoder takes DEM data as an independent channel, and the heterogeneous features extracted independently from HRSIs and DEM data are fused through channel concatenation with the CA mechanism, contrasted to the scheme in [27] which duplicates the DEM data three times to facilitate the pixel addition with the RGB channel.
@@ 3.1.1 Heterogeneous Feature Extractor
Figure 5: Architecture of Heterogeneous Feature Extractor. The structure of Coordinate Attention mechanism is indicated by the dotted line and the number on the right side of each component indicates the number of output channels
@@ 3.1.2 Multi-scale Abstract Feature Extractor
In order to distinguish relic landslides from backgrounds with high similarity, much higher abstract semantic features are required to extract from the fused features, which is implemented by MAFE. The backbone of the module is the dilated ResNet101 [7] that has been widely proven to have an excellent feature extraction capability. The ResNet101 consists of four dilated residual blocks. In order to retain more detailed features, the feature map of Block2 is concatenated with the upsampled feature map of Block4. The concatenated feature map is adaptively weighted through the SENet [18]. Then, the Atrous Spatial Pyramid Pooling (ASPP) [6] is used to effectively extract features at different scales through multiple parallel atrous convolution with different dilation rates. In addition, the global average pooling is used to gain global context information of the feature map.
@@ 3.2 Feature Extraction Enhancement Module
Figure 6: Illustration of the hyper-pixel definition. Each $8\times 8$ blue square on HRSI, DEM or Label is a hyper-pixel corresponding to a pixel of 128 channels in the feature map produced by the projection head
@@ 3.2.1 Contrastive Sample Pairs Constructor
Figure 7: Anchor Sampling schematics. On the left of Steps 1 and 2 is the feature map produced by the projection head, from which anchors are sampled. Yellow boxes and squares mark sampled landslide hyper-pixels and anchors, while green boxes and squares mark sampled background hyper-pixels and anchors. The correspondence between hyper-pixels and anchors is shown by the red dashed line
@@
In practice, the similarities between the anchor and all keys in queues are first calculated. In the queue of the same category, the top 10% $M$ keys with the lowest similarities are selected as the hard positive samples. Similarly, In the queue of the different category, the top 10% $M$ keys with the highest similarities are selected as the hard negative samples. The process of key sampling is illustrated in Fig. 4. These two types of hard samples can contribute more to the improvement of the discriminative ability [23, 36], and the remaining 90% $M$ keys are randomly sampled from the remaining keys in the queues, where $M$ is set to be 1000.
@@ 3.4.1 Reference model
We compare the HPCL-Net with the FFS-Net [27] that is considered as the baseline model. Since the FFS-Net has shown that it outperforms universal semantic segmentation models such as the UNet and Deeplabv3+ for landslide detection, we do not need to repeat the same comparison experiments. In addition, other multi-modal data fusion models [30, 14, 3, 19, 33, 43] use the data such as DSM, multi-spectral bands, aerial photograph and InSAR greatly different from our HRSIs and DEM data, where DSM and aerial photograph have much higher resolution, and InSAR contains distinctive semantic features such as deformation rate. So, it is difficult to directly migrate the HRSIs and DEM data to these models. Besides, the tasks of these models are different from ours, e.g., the landslide classification [19] and image registration [43], and the source codes of these models have not been publicly released, so we cannot reproduce their results. Considering the above cases, we have to solely compare our model with the FFS-Net.
In addition, we argue that the simple pixel-wise addition performs well in [27] because the very limited number of samples cannot support high-capacity encoder. However, when data is fully utilized in a more efficient way, a more powerful encoder is expected to enhance the feature extraction performance. Therefore, on the basis of the designed contrastive learning scheme that enables a much stronger feature extraction ability on small-sized datasets, we upgrade the encoder in [27] through channel concatenation and CA mechanism [17] when fusing heterogeneous features, so that the contrastive learning and improved encoder can bring the best in each other for efficient and reliable semantic feature extraction.

### v4

@@ Abstract
journal: Remote Sensing of Environment
[…]
[1]organization=Key Lab of Universal Wireless Communication, MOE, Beijing University of Posts and Telecommunications,city=Beijing, postcode=100876, country=China
[…]
[2]organization=China Aero Geophysical Survey and Remote Sensing Center for Natural Resources, city=Beijing, postcode=10083, country=China
[…]
[3]organization=School of Engineering and Mathematical Sciences, La Trobe University, city=Melbourne, postcode=3086, state=Victoria, country=Australia
{highlights}
Propose HPCL-Net for relic landslide detection to address the visual blur problem.
Design a dual-branch module to effectively extract and fuse heterogeneous features.
Design hyper-pixel-wise contrastive learning method for small-sized dataset problem.
Build global category queues achieving cross-image contrast to enhance effectiveness.
@@ 1 Introduction
Landslides are severe geological hazards that often results in traffic disruption, village burial, river blockage and other catastrophic incidents, posing significant threats to human life and property. For instance, Lanzhou City, China, has experienced 24 large-scale landslides since 1949, resulting in 670 fatalities and direct economic losses of 776 million RMB (Peng et al., 2019b). Therefore, accurate and efficient landslide prediction is crucial for preventing and reducing such disastrous losses. Besides, studying existing landslides could provide valuable insights for predicting potential landslides that lack fully-formed features and present the greatest risks, so this paper focuses on improving the detection of existing landslides.
@@
Traditional methods for landslide detection rely on field investigations and remote sensing data, which, though reliable, are time-consuming, inefficient, and highly dependent on experts experience, therefore hard to be widely applied. With the development of remote sensing technologies, abundant ground-observable data has been introduced for automatic landslide detection. For example, Interferometric Synthetic Aperture Radar (InSAR) data can provide deformation characteristics (Liang et al., 2023; Zhou et al., 2023), High-Resolution Satellite Image (HRSI) data can offer optical features (Lu et al., 2023; Liu et al., 2023), and Digital Elevation Model (DEM) and Digital Surface Model (DSM) data can present topographic information (Jie et al., 2018; Cao et al., 2019). There are two classic methods using HRSI: pixel-based classification and object-oriented analysis (OBA). Pixel-based methods classify pixels based on spectral features (Li et al., 2016; Zhao et al., 2017; Keyport et al., 2018) but suffer from salt-and-pepper noise and lack the ability to identify spatially continuous regions (Stumpf and Kerle, 2011; Prakash et al., 2020), whereas OBA methods improve the spatial continuity by grouping adjacent pixels into segments or objects with a homogeneity factor, leveraging spatial, texture, context, geometry, and spectral features (Blaschke, 2010; Martha et al., 2010; Lahousse et al., 2011; Hölbling et al., 2012; Pawłuszek et al., 2019). However, OBA methods require manual adjustments for thresholds such as relief and slopel, limiting their scalability.
@@
Recent machine learning methods including support vector machine, random forest, logistic regression (Chen et al., 2018a; Dou et al., 2020; Nhu et al., 2020), and convolutional neural networks (CNNs) have demonstrated improved accuracy and efficiency in landslide detection tasks. CNNs in particular are widely applied due to their excellent end-to-end self-learning ability and abstract representation capability. In Fang et al. (2020) a Siamese CNN based on dual-temporal landslide HRSI data is developed for the pixel-wise change detection of landslides and a generative adversarial network is introduced to suppress differences between two images in the time domain. In Ji et al. (2020), a 3D spatial and channel attention module is designed to enhance the extraction of landslide features. In Wang et al. (2023b), a dual path attention network is proposed to leverage both texture and spatial information in remote sensing images.
@@
Figure 1: An example to show the visual differences between an relic landslide and a new landslide. (a) relic landslide, (b) new landslide.
@@
Despite significant progress, most research focuses on detecting new landslides (Fang et al., 2020; Ghorbanzadeh et al., 2019; Ji et al., 2020; Ju et al., 2020; Cai et al., 2021; Wang et al., 2023b) rather than relic landslides (Zhang et al., 2018; Du et al., 2021; Liu et al., 2023; Lu et al., 2023). Since landslide morphology gradually changes over time, there is a significant difference in morphology between new and relic landslides. New landslides, in a state of repeated or just stopped activity, present visually obvious features, such as distinct colors and contours against background. Relic landslides, however, have been inactive for a long time and are severely affected by vegetation coverage, natural erosion, and human activities such as terracing, building houses, and constructing roads, exhibiting blurry visual features difficult to distinguish from background, as shown in Fig. 1. Therefore, relic landslides are much difficult to identify than new ones.
@@
To address this challenge, we propose a multi-modal data fusion method leveraging HRSI and DEM data, where DEM data can provide elevation features and gradient patterns that effectively complement HRSI’s optical information. The common data fusion strategies can be categorized into early fusion (Ghorbanzadeh et al., 2019; Ji et al., 2020) and late fusion (Cao et al., 2019; Peng et al., 2019a; Zeng et al., 2020; Wang et al., 2023a; Liu et al., 2023). The early fusion integrates multi-modal data at preprocessing stage such as combining several inputs from different modalities into one or mixing several datasets as a whole. For example, in Ghorbanzadeh et al. (2019), topographic images are regarded as additional channels of spectral images, and in Ji et al. (2020), satellite optical images, shape files of landslides’ boundaries and DEM data are mixed into one integral dataset. The early fusion can provide richer information and is easy to implement, but it requires powerful models, large training dataset, and massive computational resources, making it unsuitable for scenarios involving small-sized datasets. Besides, due to the fact that different modalities of data often have different semantic and syntactic features (e.g., visual features such as morphology, texture, and color from HRSIs, and terrain features like height variation from DEM), the data space after fusion is always complex and irregular, thus making further feature extraction from it quite challenging. This calls for more powerful models which place higher demands on datasets. In contrast, the late fusion independently extract heterogeneous features and then fuse them in the semantic space, which effectively obtains the essential semantic information required for the task. For example, in Cao et al. (2019), the semantic features from near infrared, red, green (IRRG) spectrum and Digital Surface Model (DSM) data are concatenated in the channel dimension; in Peng et al. (2019a), optical image features and DSM data features are contenated; in Zeng et al. (2020), two regional feature maps from optical image and Synthetic Aperture Radar data are contenated; in Liu et al. (2023), the features from HRSI and DEM data are fused by adding pixels with the same positions in spatial dimension. Therefore, considering the heterogeneity between HRSIs and DEM data, we adopt the late fsuion. Instead of pixel-wise addition (Liu et al., 2023), we fuse by channel concat since it performs better when there are significant semantic differences between heterogeneous features, such as HRSI and DEM, whereas the pixel-wise addition increases less complexity but at the expense of the discriminative ability of the model for heterogeneous features.
@@
Figure 2: Illustration of projections from a 3D landslide (bottom) into DEM (left), Contour (middle) and HRSI (right). One of the salient features of a landslide is the steep altitude descent of back and side walls, which is located in light gray areas with high elevation in DEM. The contour map is obtained by applying the Gauss-Laplace edge operator to DEM, where the back wall is parallel to the contour lines, while the side wall is vertical to the contour lines highlighted in green. The yellow arrows indicate landslide aspect. The gradient variation features in DEM are projected as the texture features in HRSI. The bottom schematic diagram of landslide is taken from Geotech (2022).
@@
In order to tackle the small-sized dataset problem, we employ contrastive learning to enhance feature extraction by distinguishing positive samples from negative ones in the feature space (Wu et al., 2018; He et al., 2020; Chen et al., 2020). Based on expert knowledge that the boundary features of back and side walls reflect the most elevation change of landslides as illustrated in Fig. 2, we propose a supervised hyper-pixel-wise contrastive learning method forcing the model to learn these local salient features in the feature space with great efficiency and reliability. The hyper pixel, defined as an area of the original image corresponding to a pixel in the high-level feature map by position, contains much rich and abundant semantic information from a local region compared to the information captured by individual pixels in pixel-wise contrastive learning.
@@ 2.1 Study area
The study area locates in the Loess Plateau at the intersection of the western end of the Western Qinling Fold Belt and the southeastern end of the Qilian Fold Belt. Specifically, the relic landslides are distributed in six counties of Gansu Province of China, i.e., Kangle, Weiyuan, Lintan, Lintao, Guanghe, and Dangchang. Fig. 3 shows the location of the study area and the landslide distribution maps from Lintao and Weiyuan. And the distribution maps of the remain four counties are presented in Appendix A(Fig. 1).
Figure 3: The geo-location of the study area and the distribution maps from Lintao and Weiyuan.
The low mountain and hilly landforms in the area are generally developed, with vertical and horizontal gullies. The stratigraphic lithologies are dominated by loose loess on the surface and weak mudstone in the lower part, with residual and alluvial parent material of soil. Besides, most of the soil moisture conditions in the area are not suitable for forest development, and the natural vegetation is grassland and shrub. Coniferous forest and deciduous broad-leaved forest are distributed in southern part of the study area since climate there is relatively humid (Zhu et al., 2023). Therefore, the vegetation coverage and the viscous soil with weak water permeability, corrosion resistance and erosion resistance are attributed to the the occurrence of landslides in the past.
@@ 2.2 Loess Plateau relic landslide dataset
The dataset used in the paper were measured in 2018 by the Gaofen-1 satellite, and contains 168 relic landslide samples cropped to the size of $512\times 512$. The HRSIs are RGB images and the DEM data are grayscale images. The images in the dataset are divided into the training, validation, and test sets in a ratio of 6:2:2. The resolution of the HRSI is 2m/pixel, and the resolution of the DEM data is originally 30m/pixel and subsequently interpolated to match the resolution of the HRSI. Since there are only 266 landslide samples, 5-fold cross validation is employed to achieve more stable experimental results, which differs from the fixed dataset partition scheme in the reference baseline model (Liu et al., 2023). Both the dataset partition scheme and the extended dataset cause the greater difference in numerical results of the baseline model.
@@ 3 Methods
Figure 4: Architecture of HPCL-Net.
@@ 3.1 Encoder
As shown in Fig. 4, the encoder consists of HFE and MAFE. Improved from our previous work (Liu et al., 2023), the newly designed encoder takes DEM data as an independent channel, and the heterogeneous features extracted independently from HRSIs and DEM data are fused through channel concatenation with the CA mechanism, contrasted to the scheme in Liu et al. (2023) which duplicates the DEM data three times to facilitate the pixel addition with the RGB channel.
@@ 3.1.1 Heterogeneous Feature Extractor
Figure 5: Architecture of Heterogeneous Feature Extractor. The structure of Coordinate Attention mechanism is indicated by the dotted line and the number on the right side of each component indicates the number of output channels.
@@ 3.1.2 Multi-scale Abstract Feature Extractor
In order to distinguish relic landslides from backgrounds with high similarity, much higher abstract semantic features are required to extract from the fused features, which is implemented by MAFE. The backbone of the module is the dilated ResNet101 (Chen et al., 2018c) that has been widely proven to have an excellent feature extraction capability. The ResNet101 consists of four dilated residual blocks. In order to retain more detailed features, the feature map of Block2 is concatenated with the upsampled feature map of Block4. The concatenated feature map is adaptively weighted through the SENet (Hu et al., 2018). Then, the Atrous Spatial Pyramid Pooling (ASPP) (Chen et al., 2018b) is used to effectively extract features at different scales through multiple parallel atrous convolution with different dilation rates. In addition, the global average pooling is used to gain global context information of the feature map.
@@ 3.2 Feature Extraction Enhancement Module
Figure 6: Illustration of the hyper-pixel definition. Each $8\times 8$ blue square on HRSI, DEM or Label is a hyper-pixel corresponding to a pixel of 128 channels in the feature map produced by the projection head.
@@ 3.2.1 Contrastive Sample Pairs Constructor
Figure 7: Anchor Sampling schematics. On the left of Steps 1 and 2 is the feature map produced by the projection head, from which anchors are sampled. Yellow boxes and squares mark sampled landslide hyper-pixels and anchors, while green boxes and squares mark sampled background hyper-pixels and anchors. The correspondence between hyper-pixels and anchors is shown by the red dashed line.
@@
In practice, the similarities between the anchor and all keys in queues are first calculated. In the queue of the same category, the top 10% $M$ keys with the lowest similarities are selected as the hard positive samples. Similarly, In the queue of the different category, the top 10% $M$ keys with the highest similarities are selected as the hard negative samples. The process of key sampling is illustrated in Fig. 4. These two types of hard samples can contribute more to the improvement of the discriminative ability (Khosla et al., 2020; Robinson et al., 2020), and the remaining 90% $M$ keys are randomly sampled from the remaining keys in the queues, where $M$ is set to be 1000.
@@ 3.4.1 Reference model
We compare the HPCL-Net with the FFS-Net (Liu et al., 2023) that is considered as the baseline model. Since the FFS-Net has shown that it outperforms universal semantic segmentation models such as the UNet and Deeplabv3+ for landslide detection, we do not need to repeat the same comparison experiments. In addition, other multi-modal data fusion models (Mondini et al., 2019; Ghorbanzadeh et al., 2019; Cao et al., 2019; Ji et al., 2020; Peng et al., 2019a; Zeng et al., 2020) use the data such as DSM, multi-spectral bands, aerial photograph and InSAR greatly different from our HRSIs and DEM data, where DSM and aerial photograph have much higher resolution, and InSAR contains distinctive semantic features such as deformation rate. So, it is difficult to directly migrate the HRSIs and DEM data to these models. Besides, the tasks of these models are different from ours, e.g., the landslide classification (Ji et al., 2020) and image registration (Zeng et al., 2020), and the source codes of these models have not been publicly released, so we cannot reproduce their results. Considering the above cases, we have to solely compare our model with the FFS-Net.
In addition, we argue that the simple pixel-wise addition performs well in Liu et al. (2023) because the very limited number of samples cannot support high-capacity encoder. However, when data is fully utilized in a more efficient way, a more powerful encoder is expected to enhance the feature extraction performance. Therefore, on the basis of the designed contrastive learning scheme that enables a much stronger feature extraction ability on small-sized datasets, we upgrade the encoder in Liu et al. (2023) through channel concatenation and CA mechanism (Hou et al., 2021) when fusing heterogeneous features, so that the contrastive learning and improved encoder can bring the best in each other for efficient and reliable semantic feature extraction.